# VPN: Visual Prompt Navigation

### v3

@@ 1 Introduction
Developing robots that can navigate based on human instructions has been a continuous research focus in artificial intelligence (AI) and robotics. Towards this goal, the research community has developed successive navigation paradigms spanning from PointGoal navigation (PointNav) (Gupta et al. 2017) and ImageGoal navigation (ImageNav) (Zhu et al. 2017) to more complex formulations including object navigation (ObjectNav) (Savva et al. 2017), vision-and-language navigation (VLN) (Anderson et al. 2018b). Despite significant progress in these visual navigation paradigms, they still have some limitations.
@@
As Figure 1 illustrates, natural language often falls short in precisely specifying navigation trajectories. This is particularly evident when describing object positions, directional transitions, or distance relationships (Chang et al. 2024). Furthermore, the pursuit of linguistic precision through exhaustive description inevitably leads to prohibitive verbosity. This exposes an inherent dilemma in human-robot interaction: linguistic precision often comes at the cost of excessive verbosity (Bonarini 2020; Wang et al. 2024a). However, overly verbose language instructions often impose high demands on users and result in poor user experiences. In contrast to language instructions, image goals offer more direct and precise navigation targets. However, they suffer from two key limitations: 1) Lack of intermediate cues. The absence of intermediate navigation cues makes it difficult for agents to plan the correct route. 2) Impractical user demands. The requirement to pre-capture a large number of goal images as various destination candidates imposes impractical demands on usability.
@@
To verify the feasibility of using visual prompts as agent navigation guidance, we explore two navigation settings: the discrete VPN setting and the continuous VPN setting (VPN-CE), transferred from the VLN benchmarks: Room-to-Room (R2R) (Anderson et al. 2018b) and R2R-CE (Krantz et al. 2020) datasets, respectively. In both settings, we design visual prompts within 2D top-view maps, where users can simply mark a sequence of key navigation waypoints, which are then connected by arrows to indicate the expected navigation trajectory. Based on such visual prompts, we construct two new datasets: R2R with visual prompts (R2R-VP) and R2R-CE with visual prompts (R2R-CE-VP), which serve as standardized benchmarks for VPN tasks in discrete and continuous environments.
To tackle the VPN task, we propose a new method called Visual Prompt Navigation Network (VPNet). In particular, we implement a discrete navigation setting of VPNet based on DUET (Chen et al. 2022b), and a continuous navigation setting based on ETPNav (An et al. 2024), where the original instruction encoders in both models are replaced with a ViT-based (Dosovitskiy et al. 2020) encoder to encode top-view maps with visual prompts. Furthermore, we introduce two data augmentation strategies to enhance navigation performance: 1) View-level augmentation, which increases view diversity by randomly altering the agent’s initial heading and rotating the visual prompt maps. 2) Trajectory-level augmentation, which increases trajectory diversity by incorporating additional trajectories from both MP3D (Chang et al. 2017) scenes and HM3D (Ramakrishnan et al. 2021) scenes. Besides, we validate how visual prompt forms, top-view map formats, and data augmentation strategies affect VPN performance.
@@ 2 Related Work
Visual Navigation requires the agent to utilize visual information to reach a specified destination. Based on the form of input instructions, it can be broadly categorized into five types: PointGoal navigation (PointNav) (Anderson et al. 2018a), object navigation (ObjNav) (Savva et al. 2017), vision-and-language navigation (VLN) (Anderson et al. 2018b), visual audio navigation (VAN) (Chen et al. 2020) and ImageGoal navigation (ImageNav) (Zhu et al. 2017). Besides, FloNa (Li et al. 2025a) leverages floor plan outlines to guide agent navigation, but it heavily relies on accurate location information of both the agent and the destination, similar to PointNav. In PointNav, the agent is provided with the relative direction and distance to the goal. In contrast, our proposed VPN guides navigation solely via visual prompts within 2D top-view maps. This design can be seen as a hybrid of VLN and ImageNav: it combines the spatiotemporal trajectory descriptions characteristic of VLN with the intuitive visual presentation style of ImageNav.
Visual Prompt (Wu et al. 2024) has recently emerged as a new paradigm that complements textual prompts by enabling more fine-grained, pixel-level instruction. Building on this idea, several studies have introduced symbolic elements, such as numbers, sketches, and arrows, to enhance natural language and improve the accuracy of Visual Question Answering (VQA) (Yang et al. 2023; Cai et al. 2024). Furthermore, recent works have explored the role of visual prompts in embodied AI (EAI). One line of work leverages generated visual prompts to improve agents’ planning capabilities, including robotic manipulation (Lee et al. 2024; Liu et al. 2024) and continuous VLN (Chen et al. 2025). However, they still rely primarily on natural language input, with visual prompts serving only a supplementary role. Another line of work, represented solely by RoVI (Li et al. 2025b), uses hand-drawn symbolic representations as the sole form of instruction for robotic manipulation. While the above work has demonstrated the role of visual prompts in EAI, using visual prompts as the sole instruction input in visual navigation remains unexplored. Unlike the relatively static scenes in robotic manipulation, the scenes in visual navigation change continuously as the agent moves. Therefore, in this work, we are the first to adopt a compact scene representation using 2D top-view maps, where visual prompts are overlaid to serve as the sole guidance for agent navigation, without relying on any textual information.
@@ 3.1 Dataset Construction
As Figure 2 illustrates, we construct the R2R-VP and R2R-CE-VP datasets through a four-step process: a) For each episode in R2R and R2R-CE, we generate a 2D top-view map of the corresponding scene from 90 MP3D (Chang et al. 2017) scenes, following a procedure similar to LED (Hahn et al. 2020). b) We then compute the pixel coordinates of all viewpoints on the top-view map and connect them sequentially with arrows to represent the agent’s navigation trajectory. c) Next, we apply a center-cropping operation to extract a square region centered on the trajectory. The side length is determined by the greater of the trajectory’s width or height, plus an additional 60-pixel margin, ensuring the full trajectory is included while maintaining spatial focus. d) Finally, to remove unnecessary black borders and tightly fit the visual prompts, we detect the bounding box of all non-zero pixels and crop accordingly. This process yields our two new datasets: R2R-VP and R2R-CE-VP. To further improve generalization and coverage, we enrich the training datasets by incorporating additional in-domain episodes from PREVALENT (Hao et al. 2020) (based on MP3D scenes) and out-of-domain episodes from ScaleVLN (Wang et al. 2023a) (based on HM3D (Ramakrishnan et al. 2021) scenes).
@@ 3.2 Dataset Analysis
Table 1 summarizes the training dataset statistics for both VLN and VPN tasks under discrete and continuous navigation settings. It includes R2R, PREVALENT (constructed from MP3D scenes), and ScaleVLN (built from both HM3D and Gibson (Xia et al. 2018) scenes). In the ScaleVLN dataset used for VPN, we retain 523 scenes from HM3D for training. In general, certain entries under VPN contain fewer scenes or episodes compared to VLN, mainly due to two reasons: 1) Regarding scenes, some scenes are excluded due to poor reconstruction quality in their corresponding top-down maps, which makes them unsuitable for annotating reliable visual prompts. For MP3D, where each scene consists of multiple floors with separate top-down views, we manually exclude floors with poor reconstruction. For the HM3D scenes, we retain only the first floor, which generally has the best reconstruction quality. We apply the above strategies to transfer the VLN datasets to visual prompt navigation, including R2R for both continuous and discrete settings, as well as ScaleVLN and PREVALENT for the discrete setting. 2) Regarding episodes, some episodes are excluded because certain viewpoints cannot be mapped to MP3D’s 3D meshes (Krantz et al. 2020) under the continuous navigation setting. As a result, only valid viewpoints from R2R-CE are retained, and any episodes with invalid viewpoints in the PREVALENT dataset are discarded.
@@ 4 VPNet
Problem formulation. In the discrete navigation setting of VPN, the environment is an undirected navigation connectivity graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ denotes navigable nodes and $\mathcal{E}$ represents the edges indicating connectivity between nodes. Starting from an initial node and guided by visual prompts, the agent is required to explore the graph $\mathcal{G}$ and navigate to the target node. The visual prompt embeddings $\mathcal{P}={\{vp_{i}\}}_{i=1}^{m}$ are encoded by a ViT-based (Dosovitskiy et al. 2020) encoder. At each time step $t$, the agent observes panoramic views $V_{t}=\{v_{t,i}\}_{i=1}^{n}$. The agent is aware of a few navigable views $\mathcal{N}(V_{t})\subseteq V_{t}$, each corresponding to a neighboring node and its spatial coordinates.
VPN in continuous environments (VPN-CE) is established over Habitat (Savva et al. 2019) simulator. At each navigation step, we employ a pretrained waypoint predictor (Hong et al. 2022) to generate navigable waypoints in continuous environments, thereby aligning the task formulation with that of VPN in discrete environments.
@@
Overview. We propose the Visual Prompt Navigation Network (VPNet) to address the newly introduced VPN task. Specifically, VPNet is based on DUET (Chen et al. 2022b) in the discrete navigation setting and ETPNav (An et al. 2024) in the continuous navigation setting. The architecture of VPNet, as illustrated in Figure 3, comprises three main components: a ViT-based Visual Prompt Encoder, a Node Embedding module, and a Graph-aware Cross-modal Encoder.
@@ 4.4 Data Augmentation
Trajectory-Level Augmentation. To enhance trajectory diversity, we introduce trajectory-level augmentation by incorporating additional navigation episodes from in-domain and out-of-domain environments. Here, in-domain refers to environments included in the R2R training set, while out-of-domain refers to those that are not. Specifically, we include in-domain trajectories from the PREVALENT dataset (Hao et al. 2020), which are based on 60 MP3D (Chang et al. 2017) scenes (Chang et al., 2017), as well as out-of-domain trajectories from the ScaleVLN dataset (Wang et al. 2023a), which are constructed using 523 HM3D (Ramakrishnan et al. 2021) scenes. This augmentation strategy enriches the training data with diverse navigation paths, thereby expanding coverage across different scene types and improving the agent’s robustness to domain shifts.
@@ 4.5 Training
Behavior cloning (BC) $\psi$ often suffers from distribution shifts between training and testing phases. To mitigate this issue, we fine-tune the VPNet using supervision from a pseudo interactive demonstrator $\psi^{*}$, following the DAgger algorithm (Ross, Gordon, and Bagnell 2011). The final loss combines Behavior cloning and DAgger supervision:
@@ 5.1 Datasets and Evaluation Metrics
We train our model using visual prompt maps corresponding to trajectories selected from the training split of R2R (Anderson et al. 2018b) and R2R-CE (Krantz et al. 2020), PREVALENT (Hao et al. 2020), and ScaleVLN (Wang et al. 2023a) datasets. For details on the validation and test splits, please refer to the appendix.
@@ 5.2 Implementation Details
Features. We adopt the ViT-B/16 (Dosovitskiy et al. 2020), pretrained on ImageNet (Deng et al. 2009), to extract the panoramic view features $W_{ij}^{v}$ for each view in panoramas and to encode the visual prompt map $\mathcal{P}_{i}$ for each navigation episode across all datasets.
@@
Seq2Seq (Anderson et al. 2018b) 6.01 53 39 7.81 28 21 7.85 27 20
RCM (Wang et al. 2019) 3.53 75 67 6.09 50 43 6.12 50 43 38
PREVALENT (Hao et al. 2020) 3.67 69 65 4.71 58 53 5.30 61 54 51
@@
HAMT (Chen et al. 2021b) 2.51 76 72 2.29 66 61 3.93 72 65 60
DUET (Chen et al. 2022b) 2.28 86 79 73 3.31 81 72 60 3.65 76 69 59
BEVBert (An et al. 2022) 2.17 88 81 74 2.81 84 75 64 3.13 81 73 62
GridMM (Wang et al. 2023b) - - - - 2.83 - 75 64 3.35 - 73 62
Lily (Lin et al. 2023) - - - - 2.90 - 74 62 3.44 - 72 60
ScaleVLN (Wang et al. 2023a) 2.12 87 81 75 2.09 88 81 70 2.27 86 80 70
NaviLLM (Zheng et al. 2024) - - - - 3.51 - 67 59 3.71 - 68 60
NavGPT-2 (Zhou et al. 2024) 1.78 88 81 75 3.13 81 72 61 3.35 78 71 60
VER (Liu, Wang, and Yang 2024) - - - - 2.80 - 76 65 2.74 - 76 66
MAGIC (Wang et al. 2024b) 1.73 89 84 80 2.22 86 79 70 2.75 82 77 69
SRDF (Wang et al. 2024d) - - - - 1.62 90 86 79 1.82 89 85 78
@@
Seq2Seq (Krantz et al. 2020) 7.12 37 35 7.37 32 30
CMTP (Chen et al. 2021a) 7.10 36 31 7.90 26 23
WPN (Krantz et al. 2021) 5.48 46 43 6.31 36 34
LAW (Raychaudhuri et al. 2021) 6.35 40 37 6.83 35 31
CM2 (Georgakis et al. 2022) 6.10 43 35 7.02 34 28
WS-MGMap (Chen et al. 2022a) 5.65 47 43 6.28 39 34
Sim2Sim (Krantz and Lee 2022) 4.67 52 44 6.07 43 36
CMA (Hong et al. 2022) 5.20 51 45 6.20 41 36
GridMM (Wang et al. 2023b) 4.21 59 51 5.11 49 41
ETPNav (An et al. 2024) 3.95 66 59 4.71 57 49
ERG (Wang et al. 2024c) 5.04 46 42 6.20 39 35
NaVILA (Cheng et al. 2024) - - - 5.22 54 49

### v4

@@ 1 Introduction
Developing robots that can navigate based on human instructions has been a continuous research focus in artificial intelligence (AI) and robotics. Towards this goal, the research community has developed successive navigation paradigms spanning from PointGoal navigation (PointNav) (gupta2017cognitive) and ImageGoal navigation (ImageNav) (zhu2017target) to more complex formulations including object navigation (ObjectNav) (savva2017minos), vision-and-language navigation (VLN) (anderson2018vision). Despite significant progress in these visual navigation paradigms, they still have some limitations.
@@
As Figure 1 illustrates, natural language often falls short in precisely specifying navigation trajectories. This is particularly evident when describing object positions, directional transitions, or distance relationships (chang2024survey). Furthermore, the pursuit of linguistic precision through exhaustive description inevitably leads to prohibitive verbosity. This exposes an inherent dilemma in human-robot interaction: linguistic precision often comes at the cost of excessive verbosity (bonarini2020communication; wang2024large). However, overly verbose language instructions often impose high demands on users and result in poor user experiences. In contrast to language instructions, image goals offer more direct and precise navigation targets. However, they suffer from two key limitations: 1) Lack of intermediate cues. The absence of intermediate navigation cues makes it difficult for agents to plan the correct route. 2) Impractical user demands. The requirement to pre-capture a large number of goal images as various destination candidates imposes impractical demands on usability.
@@
To verify the feasibility of using visual prompts as agent navigation guidance, we explore two navigation settings: the discrete VPN setting and the continuous VPN setting (VPN-CE), transferred from the VLN benchmarks: Room-to-Room (R2R) (anderson2018vision) and R2R-CE (krantz2020beyond) datasets, respectively. In both settings, we design visual prompts within 2D top-view maps, where users can simply mark a sequence of key navigation waypoints, which are then connected by arrows to indicate the expected navigation trajectory. Based on such visual prompts, we construct two new datasets: R2R with visual prompts (R2R-VP) and R2R-CE with visual prompts (R2R-CE-VP), which serve as standardized benchmarks for VPN tasks in discrete and continuous environments.
To tackle the VPN task, we propose a new method called Visual Prompt Navigation Network (VPNet). In particular, we implement a discrete navigation setting of VPNet based on DUET (chen2022think), and a continuous navigation setting based on ETPNav (an2024etpnav), where the original instruction encoders in both models are replaced with a ViT-based (dosovitskiy2020image) encoder to encode top-view maps with visual prompts. Furthermore, we introduce two data augmentation strategies to enhance navigation performance: 1) View-level augmentation, which increases view diversity by randomly altering the agent’s initial heading and rotating the visual prompt maps. 2) Trajectory-level augmentation, which increases trajectory diversity by incorporating additional trajectories from both MP3D (chang2017matterport3d) scenes and HM3D (ramakrishnan2021habitat) scenes. Besides, we validate how visual prompt forms, top-view map formats, and data augmentation strategies affect VPN performance.
@@ 2 Related Work
Visual Navigation requires the agent to utilize visual information to reach a specified destination. Based on the form of input instructions, it can be broadly categorized into five types: PointGoal navigation (PointNav) (anderson2018evaluation), object navigation (ObjNav) (savva2017minos), vision-and-language navigation (VLN) (anderson2018vision), visual audio navigation (VAN) (chen2020soundspaces) and ImageGoal navigation (ImageNav) (zhu2017target). Besides, FloNa (li2025flona) leverages floor plan outlines to guide agent navigation, but it heavily relies on accurate location information of both the agent and the destination, similar to PointNav. In PointNav, the agent is provided with the relative direction and distance to the goal. In contrast, our proposed VPN guides navigation solely via visual prompts within 2D top-view maps. This design can be seen as a hybrid of VLN and ImageNav: it combines the spatiotemporal trajectory descriptions characteristic of VLN with the intuitive visual presentation style of ImageNav.
Visual Prompt (wu2024visual) has recently emerged as a new paradigm that complements textual prompts by enabling more fine-grained, pixel-level instruction. Building on this idea, several studies have introduced symbolic elements, such as numbers, sketches, and arrows, to enhance natural language and improve the accuracy of Visual Question Answering (VQA) (yang2023dawn; cai2024vip). Furthermore, recent works have explored the role of visual prompts in embodied AI (EAI). One line of work leverages generated visual prompts to improve agents’ planning capabilities, including robotic manipulation (lee2024affordance; liu2024moka) and continuous VLN (chen2025affordances). However, they still rely primarily on natural language input, with visual prompts serving only a supplementary role. Another line of work, represented solely by RoVI (li2025robotic), uses hand-drawn symbolic representations as the sole form of instruction for robotic manipulation. While the above work has demonstrated the role of visual prompts in EAI, using visual prompts as the sole instruction input in visual navigation remains unexplored. Unlike the relatively static scenes in robotic manipulation, the scenes in visual navigation change continuously as the agent moves. Therefore, in this work, we are the first to adopt a compact scene representation using 2D top-view maps, where visual prompts are overlaid to serve as the sole guidance for agent navigation, without relying on any textual information.
@@ 3.1 Dataset Construction
As Figure 2 illustrates, we construct the R2R-VP and R2R-CE-VP datasets through a four-step process: a) For each episode in R2R and R2R-CE, we generate a 2D top-view map of the corresponding scene from 90 MP3D (chang2017matterport3d) scenes, following a procedure similar to LED (hahn2020you). b) We then compute the pixel coordinates of all viewpoints on the top-view map and connect them sequentially with arrows to represent the agent’s navigation trajectory. c) Next, we apply a center-cropping operation to extract a square region centered on the trajectory. The side length is determined by the greater of the trajectory’s width or height, plus an additional 60-pixel margin, ensuring the full trajectory is included while maintaining spatial focus. d) Finally, to remove unnecessary black borders and tightly fit the visual prompts, we detect the bounding box of all non-zero pixels and crop accordingly. This process yields our two new datasets: R2R-VP and R2R-CE-VP. To further improve generalization and coverage, we enrich the training datasets by incorporating additional in-domain episodes from PREVALENT (hao2020towards) (based on MP3D scenes) and out-of-domain episodes from ScaleVLN (wang2023scaling) (based on HM3D (ramakrishnan2021habitat) scenes).
@@ 3.2 Dataset Analysis
Table 1 summarizes the training dataset statistics for both VLN and VPN tasks under discrete and continuous navigation settings. It includes R2R, PREVALENT (constructed from MP3D scenes), and ScaleVLN (built from both HM3D and Gibson (xia2018gibson) scenes). In the ScaleVLN dataset used for VPN, we retain 523 scenes from HM3D for training. In general, certain entries under VPN contain fewer scenes or episodes compared to VLN, mainly due to two reasons: 1) Regarding scenes, some scenes are excluded due to poor reconstruction quality in their corresponding top-down maps, which makes them unsuitable for annotating reliable visual prompts. For MP3D, where each scene consists of multiple floors with separate top-down views, we manually exclude floors with poor reconstruction. For the HM3D scenes, we retain only the first floor, which generally has the best reconstruction quality. We apply the above strategies to transfer the VLN datasets to visual prompt navigation, including R2R for both continuous and discrete settings, as well as ScaleVLN and PREVALENT for the discrete setting. 2) Regarding episodes, some episodes are excluded because certain viewpoints cannot be mapped to MP3D’s 3D meshes (krantz2020beyond) under the continuous navigation setting. As a result, only valid viewpoints from R2R-CE are retained, and any episodes with invalid viewpoints in the PREVALENT dataset are discarded.
@@ 4 VPNet
Problem formulation. In the discrete navigation setting of VPN, the environment is an undirected navigation connectivity graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ denotes navigable nodes and $\mathcal{E}$ represents the edges indicating connectivity between nodes. Starting from an initial node and guided by visual prompts, the agent is required to explore the graph $\mathcal{G}$ and navigate to the target node. The visual prompt embeddings $\mathcal{P}={\{vp_{i}\}}_{i=1}^{m}$ are encoded by a ViT-based (dosovitskiy2020image) encoder. At each time step $t$, the agent observes panoramic views $V_{t}=\{v_{t,i}\}_{i=1}^{n}$. The agent is aware of a few navigable views $\mathcal{N}(V_{t})\subseteq V_{t}$, each corresponding to a neighboring node and its spatial coordinates.
VPN in continuous environments (VPN-CE) is established over Habitat (savva2019habitat) simulator. At each navigation step, we employ a pretrained waypoint predictor (hong2022bridging) to generate navigable waypoints in continuous environments, thereby aligning the task formulation with that of VPN in discrete environments.
@@
Overview. We propose the Visual Prompt Navigation Network (VPNet) to address the newly introduced VPN task. Specifically, VPNet is based on DUET (chen2022think) in the discrete navigation setting and ETPNav (an2024etpnav) in the continuous navigation setting. The architecture of VPNet, as illustrated in Figure 3, comprises three main components: a ViT-based Visual Prompt Encoder, a Node Embedding module, and a Graph-aware Cross-modal Encoder.
@@ 4.4 Data Augmentation
Trajectory-Level Augmentation. To enhance trajectory diversity, we introduce trajectory-level augmentation by incorporating additional navigation episodes from in-domain and out-of-domain environments. Here, in-domain refers to environments included in the R2R training set, while out-of-domain refers to those that are not. Specifically, we include in-domain trajectories from the PREVALENT dataset (hao2020towards), which are based on 60 MP3D (chang2017matterport3d) scenes (Chang et al., 2017), as well as out-of-domain trajectories from the ScaleVLN dataset (wang2023scaling), which are constructed using 523 HM3D (ramakrishnan2021habitat) scenes. This augmentation strategy enriches the training data with diverse navigation paths, thereby expanding coverage across different scene types and improving the agent’s robustness to domain shifts.
@@ 4.5 Training
Behavior cloning (BC) $\psi$ often suffers from distribution shifts between training and testing phases. To mitigate this issue, we fine-tune the VPNet using supervision from a pseudo interactive demonstrator $\psi^{*}$, following the DAgger algorithm (ross2011reduction). The final loss combines Behavior cloning and DAgger supervision:
@@ 5.1 Datasets and Evaluation Metrics
We train our model using visual prompt maps corresponding to trajectories selected from the training split of R2R (anderson2018vision) and R2R-CE (krantz2020beyond), PREVALENT (hao2020towards), and ScaleVLN (wang2023scaling) datasets. For details on the validation and test splits, please refer to the appendix.
@@ 5.2 Implementation Details
Features. We adopt the ViT-B/16 (dosovitskiy2020image), pretrained on ImageNet (deng2009imagenet), to extract the panoramic view features $W_{ij}^{v}$ for each view in panoramas and to encode the visual prompt map $\mathcal{P}_{i}$ for each navigation episode across all datasets.
@@
Seq2Seq (anderson2018vision) 6.01 53 39 7.81 28 21 7.85 27 20
RCM (wang2019reinforced) 3.53 75 67 6.09 50 43 6.12 50 43 38
PREVALENT (hao2020towards) 3.67 69 65 4.71 58 53 5.30 61 54 51
@@
HAMT (chen2021history) 2.51 76 72 2.29 66 61 3.93 72 65 60
DUET (chen2022think) 2.28 86 79 73 3.31 81 72 60 3.65 76 69 59
BEVBert (an2022bevbert) 2.17 88 81 74 2.81 84 75 64 3.13 81 73 62
GridMM (wang2023gridmm) - - - - 2.83 - 75 64 3.35 - 73 62
Lily (lin2023learning) - - - - 2.90 - 74 62 3.44 - 72 60
ScaleVLN (wang2023scaling) 2.12 87 81 75 2.09 88 81 70 2.27 86 80 70
NaviLLM (zheng2024towards) - - - - 3.51 - 67 59 3.71 - 68 60
NavGPT-2 (zhou2024navgpt) 1.78 88 81 75 3.13 81 72 61 3.35 78 71 60
VER (liu2024volumetric) - - - - 2.80 - 76 65 2.74 - 76 66
MAGIC (wang2024magic) 1.73 89 84 80 2.22 86 79 70 2.75 82 77 69
SRDF (wang2024bootstrapping) - - - - 1.62 90 86 79 1.82 89 85 78
DUET (R2R+PRE) 2.28 86 79 73 3.31 81 72 60 3.65 76 69 59
DUET (R2R+PRE+SCA) 2.12 87 81 75 2.09 88 81 70 2.27 86 80 70
@@
Seq2Seq (krantz2020beyond) 7.12 37 35 7.37 32 30
CMTP (chen2021topological) 7.10 36 31 7.90 26 23
WPN (krantz2021waypoint) 5.48 46 43 6.31 36 34
LAW (raychaudhuri2021language) 6.35 40 37 6.83 35 31
CM2 (georgakis2022cross) 6.10 43 35 7.02 34 28
WS-MGMap (chen2022weakly) 5.65 47 43 6.28 39 34
Sim2Sim (krantz2022sim) 4.67 52 44 6.07 43 36
CMA (hong2022bridging) 5.20 51 45 6.20 41 36
GridMM (wang2023gridmm) 4.21 59 51 5.11 49 41
ETPNav (an2024etpnav) 3.95 66 59 4.71 57 49
ERG (wang2024graph) 5.04 46 42 6.20 39 35
NaVILA (cheng2024navila) - - - 5.22 54 49
ETPNav (R2R+PRE) 3.95 66 59 4.71 57 49